2l+2w=228 L=w+42 How do you find the length(L) and width(w)?

Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

2l+2w=228 L=w+42 How do you find the length(L) and width(w)?
-----
Substitute for "L" and solve for "W":
2(W+42) + 2W = 228
W+42 + W = 114
2W = 72
Width = 36 units
----
Solve for "L"
L = W + 42
L = 78 units
==============
Cheers,
Stan H.
================

Answer by Cromlix(4381)

2L + 2W = 228
L = W + 42
Substitute (w + 42) for L
2(W + 42) + 2W = 228
2W + 84 + 2W = 228
4W = 228 - 84
4W = 144
W = 36
Therefore Width = 36 and Length = (W + 42)
= 36 + 42 = 78
Hope this helps
:-)